### v2

# thanks: These authors contributed equallythanks: These authors contributed equally

Abstract
Computing excitonic spectra in quasicrystal and super-moiré systems constitutes a formidable challenge due to the exceptional size of the excitonic Hilbert space. Here, we demonstrate a tensor-network method for the real-space Bethe-Salpeter Hamiltonian, allowing us to access the spectra of an excitonic $10^{18}$-dimensional Hamiltonian, and enabling the direct computation of bound-exciton spectral functions for systems exceeding one billion lattice sites, several orders of magnitude beyond the capabilities of conventional approaches. Our method combines a tensor-network encoding of the real-space Bethe-Salpeter Hamiltonian with a Chebyshev tensor network algorithm. This strategy bypasses explicit storage of the Hamiltonian while preserving full real-space resolution across widely different length scales. We demonstrate our methodology for one- and two-dimensional super-moiré systems, achieving the simultaneous resolution of atomistic and mesoscopic structures in the excitonic spectra in billion-size systems, showing exciton miniband formation and moiré-induced spatial confinement. Our results establish a real-space methodology enabling the simulation of excitonic physics in large-scale quasicrystal and super-moiré quantum matter.
Introduction
Two-dimensional (2D) van der Waals materials provide an excellent platform for engineering excitonic matter owing to their enhanced Coulomb interactions. This has led to pronounced excitonic features that remain stable even at room temperature [28, 75, 48], making these platforms ideal for technological applications [87, 42, 58, 64, 38, 2]. Importantly, control over twist angle and lattice mismatch in multilayer transition metal dichalcogenides (TMD) has enabled the realization of moiré excitons [88, 67, 65, 59, 17, 81, 19] by producing an effective potential landscape that modulates the coupling of electron and hole pairs across both layers, leading to emergent phenomena, such as periodic confinement and excitonic minibands. These excitons can also become strongly localized at moiré trapping sites, forming programmable arrays of quantum emitters or allowing the engineering of exciton lattices that function as artificial quantum simulators [88, 29, 19].
Microscopically, excitonic phenomena can be predicted from the Bethe–Salpeter equation (BSE), which incorporates the single-particle electronic structure and the screened electron–hole interaction [23, 79, 37, 14, 44, 41, 8]. While for periodic systems a momentum space formulation of excitons is successful, for incommensurate modulations, such as those arising in moiré materials, a fully real-space formulation for excitonic effects becomes inevitably necessary. In these cases, conventional methods face severe computational limitations, as even at the single-particle level, real-space methods scale poorly with system size [79, 84, 30, 77, 41, 8]. For excitons, the situation is more demanding, since the two-particle Hilbert space scales quadratically with the number of lattice sites. In super-moiré materials, where characteristic length scales can exceed those of standard moiré systems by orders of magnitude, the effective lattice may involve millions or even billions of sites [12, 90], rendering brute-force approaches computationally prohibitive.
In the single-particle case, quantum many-body solvers based on tensor networks (TNs) have enabled the treatment of exponentially large Hamiltonians [24, 72, 3, 47]. These approaches avoid explicit storage of large matrices by representing operators and states with TNs: many-body factorizations that permit algebraic manipulations in exponentially large Hilbert spaces [82, 15, 5, 53, 66, 71, 22, 21, 89, 31, 54, 11]. The required tensor representations are constructed using tensor cross interpolation techniques [56, 55, 62, 34, 52, 60, 61]. Similarly, the employment of TNs to encode classical functions has led to speedups of several orders of magnitude across diverse computational settings [78, 69, 20, 51, 73, 34, 35, 63], including applications in computational chemistry [36] and dynamical simulations [57, 27, 49, 9, 13, 16]. These developments have enabled, among others, the study of quasicrystalline mosaics of topological Chern states [3], the self-consistent solution of interacting super-moiré systems at scales exceeding one billion lattice sites [72], and the direct computation of spatially-dependent momentum-resolved spectral function of large nonperiodic systems [47].
Here, we demonstrate a tensor-network (TN) based methodology to access the real-space exciton spectra in super-moire systems. Our methodology relies on combining a real-space TN representation of the electron–hole Hilbert space, and tensor-network kernel polynomial Chebyshev algorithms with a high-resolution spectral kernel. Our approach enables the solution of the real-space BSE on lattices exceeding one billion sites, both in 1D and 2D, surpassing by many orders of magnitude the capabilities of conventional BSE solvers [41, 8]. We show that these techniques allow probing the emergence of exciton moire minibands, together with imaging the microscopic exciton localization, both in quasicrystalline and super-moire systems.
Tensor Network Methodology
For concreteness, we formulate our algorithm in a minimal two-band model that captures excitonic physics, and we solve it with our TN methodology. We take an interacting Wannier Hamiltonian for the valence and conduction band of a semiconductor as $\hat{H}=\hat{H}_{c}+\hat{H}_{v}+\hat{H}_{U}$, with
$\begin{split}\hat{H}_{b}&=\sum_{l,m}t_{lm}c_{b,l}^{\dagger}c_{b,m}+\sum_{l}V_{b,l}c_{b,l}^{\dagger}c_{b,l},\\ \hat{H}_{U}&=\sum_{l}U_{l}c_{c,l}^{\dagger}c_{c,l}c_{v,l}^{\dagger}c_{v,l},\end{split}$ (1)
where $c^{\dagger}_{c,l}$ ($c^{\dagger}_{v,l}$) creates an electron in the Wannier conduction (valence) band at site $l$. The band index is denoted by $b=c,v$ for conduction and valence, respectively. The hopping amplitudes are given by $t_{lm}$, while $V_{c,l}=-V_{v,l}$ represents the on-site potential, and $U_{l}$ describes the local screened electrostatic interaction between a conduction-band Wannier state and a valence-band Wannier state. This Hamiltonian stems from projecting a microscopic interacting electron model onto the low-energy subspace spanned by the top valence and bottom conduction bands, performing a Wannier transformation to real space, and retaining the local Coulomb interaction between valence and conduction bands.
Figure 1: (a) Schematic of the exciton Hamiltonian, with purple corresponding to the conduction band and light brown to the valence band, together with their corresponding single-particle basis. (b) The encoding of the two-particle $2^{2L}\times 2^{2L}$ Hamiltonian matrix into a $2L$ pseudo-spin MPO. (c) The combination of the single-particle Hamiltonians and the interaction kernel using interleaved ordering.
[…]
$\hat{H}_{X}=\hat{T_{c}}\otimes\mathbbm{1}-\mathbbm{1}\otimes\hat{T_{v}}-\hat{U}$, (2)
where $\hat{T}_{c}$ contains both the conduction hopping term and an additional contribution from the local Coulomb interaction, $\hat{T}_{v}$ is the contribution arising from the valence band, and $\hat{U}$ couples electrons and holes on the same site. Its matrix elements are $U_{klmn}=U_{k}\delta_{kl}\delta_{mn}\delta_{ln}$ and, as such, cannot be written as a product operator (see SM [1] for details).
The key strategy of our methodology lies in the encoding of $2^{L}$ real-space sites of the Wannier tight-binding Hamiltonian into $L$ pseudo-spin indices [24, 72, 3, 47] for each sector of the BSE Hamiltonian, effectively resulting in a system of $2L$ pseudo-spin degrees of freedom. In this pseudo-spin basis, the single-particle Hamiltonian of each sector can be written as a TN matrix product operator(MPO) of the form $\mathcal{H}_{(s_{1},s_{2},...,s_{L}),(s^{\prime}_{1},s^{\prime}_{2},...,s^{\prime}_{L})}={\Gamma}^{s_{1},s_{1}^{\prime}}_{1}{\Gamma}^{s_{2},s_{2}^{\prime}}_{2}\cdots{\Gamma}^{s_{L},s_{L}^{\prime}}_{L}$. Each tensor $\Gamma_{r}$ carries four indices: two virtual indices, which are contracted with neighboring tensors, and two physical indices associated with the local 2D Hilbert space $s_{r},s^{\prime}_{r}$. The virtual indices have dimension $\chi_{r}$, referred to as the bond dimension, which determines the expressiveness of the local tensor. The explicit form of the tensors can be obtained with quantics tensor cross interpolation [62, 78, 69, 20, 51, 73, 34, 35], enabling the representation of spatially varying Hamiltonians as a TN. Considering the first two terms of Eq. 2, a straightforward many-body encoding consists of two independent pseudo-spin chains of length $L$, which together form a composite chain of length $2L$, reflecting the direct-product structure of the corresponding operators. However, in this pseudo-spin representation, the interaction operator is expressed as $\hat{\mathcal{U}}=\prod_{r=1}^{L}\left(\frac{1}{2}\mathbbm{1}_{r}\otimes\mathbbm{1}_{r+L}+2\sigma^{z}_{r}\sigma^{z}_{r+L}\right)$. This specific form couples the two pseudo-spin chains by sequentially ordering all electron sites before all hole sites. Such a layout leads to an exponential growth of the bond dimension at the interface between the two MPOs as the system size increases. To mitigate this, we use a more efficient interleaved ordering [52, 78] of the sites, which maintains a constant bond dimension for the interaction operator regardless of the system size. In essence, rather than linking the individual pseudospin blocks with a single internal bond at their interface, we adopt an interleaved ordering in which the tensor cores are rearranged so that each local electron pseudospin degree of freedom is directly coupled to its corresponding hole pseudospin degree of freedom. The MPO construction of the Hamiltonian is illustrated in Fig. 1  [^1] The technical details of the implementation of the interleaved BSE Hamiltonian is discussed in the SM [1].
[…]
In practice, the expansion is truncated to $N_{\mu}$ terms. Within the KPM framework, the inclusion of the Jackson kernel [33] yields an effective energy resolution of order $1/N_{\mu}$, thereby setting the scale of the associated spectral broadening [80]. For 2D systems, however, the rapid growth of the bond dimension with the number of Chebyshev moments necessitates an alternative strategy. To keep $N_{\mu}$ moderate while maintaining spectral resolution, we employ a high-order delta-Chebyshev (HODC) kernel [86], which enables independent control of the KPM truncation error through additional parameters $\eta$ and $m$, achieving a faster convergence of order $\eta^{m}\sim 1/N_{\mu}^{m}$(see SM [1] for details).
We now demonstrate the capabilities of our TN framework by applying it to two representative examples that require the simultaneous resolution of both macroscopic and atomistic length scales.
1D incommensurate super-moiré potential
Figure 2: (a) Spectral band in real space of the bound super-moiré excitons, together with a zoom of the central region of linear extent $\ell=64$. (b) Super-moiré on-site potential modulation, with a corresponding zoom of the central region of length $\ell=64$. The outlined region is not shown to scale with respect to the full domain and serves only as a visual guide.
The first example is a 1D system subject to a modulated single-particle on-site potential of the form
$\begin{split}V_{i}&=V_{0}\left[1+0.1\cos\left(\frac{\pi x_{i}}{3\sqrt{5}}\right)+0.1\cos\left(\frac{5\pi x_{i}}{N\sqrt{3}}\right)\right],\end{split}$ (4)
where the two modulations are incommensurate with each other. This model captures moiré physics in quasi-one-dimensional van der Waals materials [32, 43, 4] or multi-walled nanotubes [7, 50, 74, 26, 76]. The first cosine induces a small-scale modulation, while the second induces a large-scale one. We also consider nearest-neighbor uniform hopping $t$, a constant electron-hole interaction strength $U=5t$, and an on-site potential strength $V_{0}=3t$. Figure 2(a) shows the large-scale modulation Eq. 4 for a system of $N=2^{30}$ sites and Fig. 2(b) the corresponding large-scale LDOS as a function of energy. The isolated exciton band in real-space closely follows the spatial profile of the on-site modulation potential. In Fig. 2(c) and Fig. 2(d) provide a zoom into the central region of length $\ell=64$ of the same quantities as in Fig. 2(a,b). The spatial modulation and its imprint in the LDOS are also well resolved at the atomic scale. Furthermore, the incommensurability of the small-scale modulation frequency gives rise to miniband splitting, which is visible at both length scales through the modulated small gap.
2D quasiperiodic incommensurate super-moiré potential
Figure 3: (a) 2D exciton LDOS $\rho(x,y,E=E_{X})$ evaluated at the bound-exciton energy $E=E_{X}$, together with a zoom of the central region of area $A=128\times 128$. (b) Corresponding on-site potential, with an analogous zoom of the central region of area $A=128\times 128$. The square outline is not drawn to scale relative to the full domain and is included as a visual guide.
We now demonstrate the methodology on a 2D super-moiré quasicrystal. We consider a square lattice with an on-site potential exhibiting an eight-fold quasiperiodic, incommensurate modulation at both small and large scales.
$V(\mathbf{r})=V_{0}\left[1+\sum_{n=1}^{4}\left(\Delta_{\alpha}\cos\left(\alpha\mathbf{k}_{n}\cdot\mathbf{r}\right)+\Delta_{\beta}\cos\left(\beta\mathbf{k}_{n}\cdot\mathbf{r}\right)\right)\right],$ (5)
where $\Delta_{\alpha}$ sets the modulation amplitude at the atomic scale, while $\Delta_{\beta}$ determines the super-moiré modulation amplitude. The eightfold pattern is created by the superposition of four wave vectors given by $\mathbf{k_{n}}=R^{n}(\pi/4)\left[2\pi,0\right]^{T}$, with $R(\pi/4)$ denoting the 2D rotation matrix corresponding to an angle of $\pi/4$ radians. This potential can result from the superposition of two square lattice monolayers rotated at $45^{\circ}$ relative to each other, as in twisted GeX/SnX moiré superlattices [85]. We set $\Delta_{\alpha}=0.15$, $\Delta_{\beta}=0.05$, and take moiré length scales $\alpha=2/(5\sqrt{5})$ and $\beta=16/(\sqrt{3}N_{x})$. The uniform nearest-neighbor hopping is $t$, the on-site modulation potential amplitude $V_{0}=6t$, and the local Coulomb interaction $U=12t$. The large-scale modulation landscape of the on-site potential is shown in Fig. 3(a) for a system of size $N_{x}=N_{y}=2^{15}$, corresponding to more than a billion lattice sites. In Fig. 3(b), the corresponding LDOS at the bound exciton energy $E_{X}=-4t$ is depicted, where one can observe that the spectral weight peaks near the minima of the on-site potential, indicating spatial confinement of the excitons in a pattern with eight-fold rotational symmetry. In Fig. 3(c) and Fig. 3(d), a zoom into a central region of area $A=128\times 128$ is shown for the on-site potential and LDOS respectively. Similarly to the super-moiré length scale, the same eightfold-symmetric confinement pattern at the minima of the potential can be observed at the atomic length scale.
Discussion
The present results establish a scalable real-space framework for solving the BSE in regimes that are inaccessible to conventional approaches. In contrast to standard implementations, whose cost scales quadratically with the number of lattice sites and which require explicit storage of the two-particle Hamiltonian, our formulation exploits the TN structure to achieve logarithmic scaling with system size and efficient TN compression. This enables the treatment of super-moiré systems with effective Hamiltonian dimensions of order $10^{18}$, a regime far beyond existing methodologies. It is worth noting that several limitations remain. In 2D, the rapid growth of bond dimensions with the number of Chebyshev moments places us near the threshold of efficient compressibility, currently limiting our accuracy to approximately forty moments. While the HODC kernel partially mitigates this by resolving spatial features of the LDOS at lower expansion orders, alternative architectures such as tree TNs [68] may be better suited to handle 2D correlations and further enhance performance. While we have focused here on local electron–hole interactions in minimal 1D and 2D lattices for concreteness, the framework can be systematically extended to incorporate long-range Coulomb interactions and exchange terms within the same TN formalism, as well as to more complicated lattice geometries. Such extensions would facilitate the quantitative modeling of realistic moiré heterostructures while preserving their real-space scalability demonstrated here. It is finally worth noting that the inclusion of long-range interactions and exchange terms may potentially accelerate the growth of the bond dimension, further motivating the development of alternative TN architectures. Beyond the specific exciton problem we have focused on here, our methodology provides a general strategy for addressing two-body correlation problems directly on real-space lattices [45]. This technique can be applicable to other correlated states, including the formation of bipolaron states in moire systems [10, 46, 18] and bimagnon bound states in moire magnetic insulators [70, 83, 40, 25, 39, 6]. In all such cases, under appropriate circumstances, computations involve a two-particle Hilbert space whose direct treatment becomes prohibitive at large system sizes, yet can be efficiently encoded within the present TN framework.
Conclusion
We have demonstrated a TN methodology that enables direct real-space calculations of bound-exciton spectral functions in exponentially large lattices. Our approach combines a pseudo-spin TN encoding of the BSE Hamiltonian, including both the single particle and many-body interaction in the full two-particle excitonic space. Our algorithm leverages an interleaved representation of the electron–hole BSE space, together with a TN-based kernel polynomial expansion enhanced by a higher-order delta Chebyshev spectral kernel. We demonstrated our methodology for 1D and 2D incommensurate super-moiré potentials with more than a billion sites, demonstrating the simultaneous resolution of atomistic and mesoscopic structures in the exciton LDOS. In both cases, we show the bound exciton spectral signatures in real space following the spatial modulation landscape. In 1D, we observed exciton miniband formation in real space and in 2D spatial confinement patterns at a specific exciton energy. Our approach provides a scalable real-space framework for exciton physics in moiré and super-moiré matter, offering a technique enabling the study of excitonic phenomena in super-moire materials.
Acknowledgements.
Acknowledgments We acknowledge the computational resources provided by the Aalto Science-IT project and the financial support from InstituteQ, the Research Council of Finland (No. 370912), the Finnish Ministry of Education and Culture through the Quantum Doctoral Education Pilot Program (QDOC VN/3137/2024-OKM-4), the Finnish Quantum Flagship (No. 358877, Aalto University), the Finnish Centre of Excellence in Quantum Materials QMAT (No. 374166), and the ERC Consolidator Grant ULTRATWISTROICS (No. 101170477). LE acknowledges the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW). We thank P. Liljeroth, S. Kezilebieke, P. Shen, E. Castro, B. Amorim, X. Waintal, A. Akhmerov, and C. M. Smith for useful discussions. The code used for this work can be consulted at [anouarrepo].